# Elastic scattering at $\sqrt{s}=6$ GeV up to $\sqrt{s}=13$ TeV

### v1

@@ V.1 Electromagnetic amplitudes and phase factor
The electromagnetic amplitude can be calculated in the framework of QED in the one-photon approximation,
@@
$\Phi^{em}_{2}(t)=\alpha\frac{f_{2}^{2}(t)}{4m^{2}},$
$\Phi^{em}_{4}(t)=-\phi^{em}_{2}(t),\Phi^{em}_{5}(t)=\alpha\frac{s}{2m\sqrt{|t|}}f_{1}^{2}.$ (33)
For numerical calculations let us take
$-t=x(1)$ momentum transfer ( $x(1)$ is taken from experimental data) and $q^{2}=-t$ GeV2.
( $fc1=1$ for $pp$ and $fc1=-1.$ for $p\bar{p}$ reactions),
$\Phi_{1}^{em}=-fc1\ \alpha_{em}(f1t^{2}/q^{2})e^{i\alpha_{em}\varphi_{CN}},$
$\Phi_{2}^{em}=fc1\alpha_{em}f2t^{2}/(pm24)e^{i\alpha_{em}*\varphi_{CN}},$
$\Phi_{3}^{em}=-fc1\alpha_{em}f1t^{2}/q^{2}e^{i\alpha_{em}*\varphi_{CN}},$
$\Phi_{4}^{em}=-\Phi_{2em},$
$\Phi_{5}^{em}=-fc1\alpha_{em}f1tf2t/(2m_{p}q)e^{i\alpha_{em}\varphi_{CN}}.$
The Coulomb-hadron interference phase is calculated with the dipole electromagnetic form factor [78]. For calculation of the Coulomb-hadron phase we take the energy dependence of the slope in form $b_{sl}=6+0.75\ln(s)$.
@@
$\phi_{b}=(\Lambda^{2}+q^{2})^{2}/(\Lambda^{4}(4.\Lambda^{2}+q^{2})^{2}*q\sqrt{(}4\Lambda^{2}+q^{2}));$
@@
$\phi_{CN}=-\ln{b_{sl}q^{2}/2.}+\gamma+\ln(1.+8.d0/(b_{sl}al2))-\phi_{1-3})$
.                  Refer to caption
Figure 4: The comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=13$ TeV, $p\bar{p}$ scattering at $\sqrt{s}=53$ GeV and $pn$ scattering at $\sqrt{s}=23$ GeV
.                 Refer to caption
Figure 5: The HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=13$ TeV.
@@ V.2 The hadron scattering amplitude
.          Refer to caption
Figure 6: The comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=30.6$ GeV (solid line and full ) and $\bar{p}p$ at $\sqrt{s}=30.4$ GeV (dashed line and open triangles). In both cases the additional normalization $n=1$). scattering.
Figure 8: [top] $sigma_{tot}(s)$: dashed line - $p\bar{p}$ scattering, solid line -$pp$ scattering; [middle] $\rho(s,t=0)$: dashed line - $p\bar{p}$ scattering, solid line -$pp$ scattering; [down] the real part of the scattering amplitude: dashed line - $p\bar{p}$ scattering, solid line -$pp$ scattering.
@@
At small $t$,the CNI region, there are two contributions coming from the electromagnetic and the strong interaction.
On the one hand, the interference of such contribution gives the possibility to determine the size of the real part of the scattering amplitude. On the other hand, to determine the form of the imaginary part of the hadronic amplitude, it is necessary to extract the Coulomb contribution and the interference term. As the electromagnetic amplitude and its main contribution at $t\rightarrow 0$ are well known, the measure of experimental data at very small $t$ gives the possibility to improve the normalization of the differential cross sections.
The existence of the different sets of the experimental data also allows to improve the separate normalization of the experimental data. To compare the different sets it is need to have some gauge.
Let us take the calculations of the differential cross sections carried out in the framework of the new High energy generalized structure (HEGS) model [38, 39] as such a gauge. The model has only a few free parameters and it quantitatively describes experimental data in a wide domain of the momentum transfer, including the data in the CNI region, in a very wide energy region (from $\sqrt{(}s)=6$ GeV up to LHC energies simultaneously with the same numbering of the free parameters. The HEGS model assumes a Born term for the scattering amplitude which gets unitarization via the standard eikonal representation to obtain the full elastic scattering amplitude. The scattering amplitude has exact $s\leftrightarrow u$ crossing symmetry as it is written in terms of the complexified Mandelstam variable $\hat{s}=se^{-i\pi/2}$ that determines its real part. The scattering amplitude also satisfies the integral dispersion relation at large $s$. It can be thought of as the simplest unified analytic function of its kinematic variables connecting different reaction channels without additional terms for separate regions of momentum transfer or energy. Note that the model reproduces the diffraction minimum of the differential cross section in a wide energy region [69]. HEGS model describes the experimental data at low momentum transfer, including the Coulomb-hadron interference region, and hence includes all five electromagnetic spin amplitudes and the Coulomb-hadron interference phase.
@@
$F_{Pom2}^{Born}(s,t)=$ $h_{Pom2}\ G^{2}_{em}(t)F_{a}(s,t);$ (35)
$F_{Pom3}^{Born}(s,t)=$ $h_{Pom3}\ A^{2}_{gr}(t)F_{b}(s,t);$
$F_{Odd3}^{Born}(s,t)=$ $h_{Odd3}\ A^{2}_{gr}(t)F_{c}(s,t);$
where $F_{a}(s,t)$ and $F_{b,c}(s,t)$ have the standard Regge form:
$F_{a}(s,t)\ =\hat{s}^{\epsilon_{1}}\ e^{B(\hat{s})\ t};\ \ \ F_{b,c}(s,t)\ =\hat{s}^{\epsilon_{1}}\ e^{B(\hat{s})/4\ t},$ (36)
with $\hat{s}=s\ e^{-i\pi/2}/s_{0}$ ; $s_{0}=4m^{2}_{p}\ {\rm GeV^{2}}$, and $h_{Odd3}=ih_{3}t/(1-r_{0}^{2}t)$. The intercept of all main parts of the scattering amplitudes $1.111+\Delta_{1}=$ 1.11 was chosen as arithmetic means from different works including inelastic scattering. Hence, at the asymptotic energy we have the universality of the energy behavior of the elastic hadron scattering amplitudes. The main part of the slope of the scattering amplitude has the standard logarithmic dependence on the energy $B(s)=\alpha^{\prime}\ ln(\hat{s})$ with $\alpha^{\prime}=0.24$ GeV-2. It is taken with some correction (eq. 16).
@@
where both (electromagnetic and gravitomagnetic) form factors are used. The constant parameters are determined by the fitting procedure.
So, to extend the model to low energies, it is necessary to take into account the contributions of the second Reggions. To avoid a substantially increasing number of fitting parameters we introduce the effective terms which represent the contributions of the sums of different Reggions. For the terms with the energy dependence order $1/\sqrt{\hat{s}}$ we take for the $pp$ scattering the cross even part in the form
$F_{R1}=ih_{R1}/\sqrt{\hat{s}}e^{b_{R1}t\ln{\hat{s}}};$ (38)
and for the $p\bar{p}$ the cross odd term in the form
$F_{R1}=h_{R1}/\sqrt{\hat{s}}e^{b_{R1}t\ln{\hat{s}}}$ (39)
where the value $b_{R1}$ is fixed by unity. For the cross even part fast decreasing with growing energy, we take term in the form
$F_{cr.ev.}=h_{R2}/\hat{s}e^{b_{R2}t\ln{\hat{s}}};$ (40)
The model is very simple from the viewpoint of the number of terms of the scattering amplitude and fitting parameters. There are no any artificial functions or any cuts which bound the separate parts of the amplitude by some region of momentum transfer. In the framework of the model, the description of experimental data was obtained simultaneously at the large momentum transfer and in the Coulomb-hadron region in the energy region from $\sqrt{s}=6$ GeV up to LHC energies. The model gives a very good quantitative description of the recent experimental data at $\sqrt{s}=13$ TeV [75].
The final elastic hadron scattering amplitude is obtained after unitarization of the Born term. So, at first, we have to calculate the eikonal phase
@@
An example of such description is shown in Fig. 6 for proton-proton at $\sqrt{s}=30.6$ GeV and for proton-antiproton scattering at $\sqrt{s}=30.4$ GeV. The difference comes from the interference term of the Coulomb and the real part of the hadron amplitudes. The $\chi^{2}/N=1.1$ for $pp$ and $\chi^{2}/N=1.03$ for $p\bar{p}$ scattering with additional normalization $k_{pp}=1.005$ and $k_{p\bar{p}}=0.974$, respectively. The figure presents the data without additional normalization. Note that this $\chi^{2}$ was obtained as separate part of the full fit of the $4327$ experimental data in a wide energy and momentum transfer region [39].
Let us compare the predictions of the HEGS model for the differential elastic cross section at small $t$ with the LHC data. In the fitting procedure only the statistical errors are taken into account. The systematic errors are reflected through an additional normalization coefficient which is the same for all the data of a given set. The different normalization coefficients have practically random distributions at small $t$.
@@
The data of the TOTEM at $\sqrt{s}=7$ TeV are consistent and their mean value is equal to $98.5$ mb. The ATLAS Collaboration, using their differential cross section data in a region of $t$ where the Coulomb-hadron interference is negligible, obtained the value $\sigma_{tot}=95.35\pm 2.0$ mb. The difference between the two results, $\sigma_{tot}(s)$(T.) - $\sigma_{tot}(s)$(A.) = $3.15$ mb, is about 1 $\sigma$. At $\sqrt{s}=8$ TeV, the measured value of $\sigma_{tot}$ grows, especially in the case of the TOTEM Collaboration and the difference between the results of the two collaborations grows to $\Delta(\sigma_{tot}(s)$(T.) - $\sigma_{tot}(s)$(A.) $=5.6$ mb, i.e. 1.9 $\sigma$. This is reminiscent of the old situation with the measurement of the total cross sections at the Tevatron at $\sqrt{s}=1.8$ TeV via the luminosity-independent method by different collaborations.
Of course, we cannot say that the normalization of the ATLAS data is better than that of the TOTEM data simply because it coincides with the HEGS predictions. But this exercise may point to the main reason for the different values of the total cross sections obtained by the two collaborations. This does not exclude some further problems with the analysis of the experimental data, e.g. those related to the analysis of the TOTEM data at $\sqrt{s}=7$ TeV [68].
The position of the diffraction minimum $t_{min}(s,t)$ moves to low momentum transfer continuously [41]. It is interesting that the velocity of changing the position of the diffraction minimum changes very slowly. For example, from ISR energy $\sqrt{s}=53$ GeV up to SPS energy $\sqrt{s}=540$ GeV such position changes with a speed of $0.11$ GeV2 per $100$ GeV. Between $\sqrt{s}=540$ GeV and $\sqrt{s}=7$ TeV such speed is two times less and equals $0.006$, at last between $7$ and $13$ TeV the position of minimum changes with a speed of $0.002$ per $100$ GeV. Approximately, scaling of this process can be represented as $t_{min}\ln{s/s_{0}}=const$. After the second bump the slope of the differential cross sections increases with energy. It corresponds to the grows of the slope of the diffraction peak.
@@
The behavior of the imaginary part of the scattering amplitude over momentum transfer is presented in Fig. 7 for the energies $\sqrt{s}=0.51,7,13$ TeV. Again, we can see the point of crossover in the region of $|t|=0.2$ GeV2. Despite the essential grows of the size of the imaginary part of the scattering amplitude at very small momentum transfer, its slope slightly changes with $t$ in the region of the Coulomb nuclear interference. The size of slope is practically proportional to the size of the total cross section in that region. However at larger $t$, for example at $|t|=0.1$ GeV2, it grows essentially faster.
@@
It should be noted that the size of the slope of the differential cross sections is determined in that region of $t$ by the CNI interference term which is proportional to $\alpha/t$ It allows us to analyze [71] the first points of the unique experiment carried out by the ATLAS Collaboration [31]. The point of $t$, where the imaginary part changes its sign, determines the position of the diffraction minimum. But it slightly moves at some large $t$ by the contribution of the real part of the elastic hadron scattering amplitude.
Figure 11: The size of the $\rho(s,t)$ ratio of the real to imaginary part of the hadron scattering amplitude is calculated in the model at the energy $\sqrt{s}=0.51,7,13$ TeV depending on $s$ and $t$.
It is interesting that the form of the real part of the hadron elastic scattering amplitude it is similar to its imaginary part. Of course, they are not proportional to each other as their connection has to satisfy the dispersion relations [58] which require, for example, the changing size of the real part at sufficiently small momentum transfer. Really, in Fig. 11, we see that the real part change its sign essentially earlier than the position of the diffraction minimum. At LHC energies this happens in the area of momentum transfer approximately equal to $0.2$ GeV2. But before that, we can see the crossing point at $|t|\sim 0.06$ GeV2. Likely, the behavior of the imaginary part the slope of the real part at very small momentum transfer is also practically proportional to the size of the total cross section at different energies and grows at larger momentum transfers. It can be see that the real part at LHC energies has the negative maximum at approximately $|t|=0.3$ GeV2 situated near the diffraction minimum. Hence, it essentially impacts the form and size of the diffraction minimum in the differential cross sections. In Fig. 11, the ratio $\rho(s,t)$ of the real to imaginary part of the hadron elastic scattering amplitude is presented for different energies. Such a complicated structure of $\rho(s,t)$ are determined by the changes of the sign of the real and imaginary parts of the scattering amplitude. At small momentum transfer, the size of $\rho(s,t)$ is small as the real part changes its sign. Contrary, when the imaginary part changes its sign, the size of $\rho(s,t)$ grows very faster. The energy dependence of $\rho(s,t)$ is due to the movement of the position of the diffraction minimum, hence with the energy dependence of the imaginary part of the scattering amplitude.
Especially the size and energy dependence of the real part of the hadronic amplitude impact on the differential cross sections at CNI region. The model descriptions of differential cross section at small momentum transfer and at low energies are presented in Figures 5, 15-17. In these figures are presented the experimental data with high precision and we show only statistical errors, which were taking into account in out fitting procedure. Obviously, that the model reproduce the experimental data very good at the wide energy region.

### v3

@@ V.1 Electromagnetic amplitudes and phase factor
The electromagnetic amplitude can be calculated in the framework of QED in the one-photon approximation
@@
$\Phi^{em}_{2}(t)=\alpha\frac{f_{2}^{2}(t)}{4m^{2}},\Phi^{em}_{4}(t)=-\phi^{em}_{2}(t),$
$\Phi^{em}_{5}(t)=\alpha\frac{s}{2m\sqrt{|t|}}f_{1}^{2}.$ (31)
The Coulomb-hadron interference phase is calculated with the dipole electromagnetic form factor [79]. For calculation of the Coulomb-hadron phase we take the energy dependence of the slope in the form $b_{sl}=6+0.75\ln(s)$.
@@
$\phi_{b}=(\Lambda^{2}+q^{2})^{2}/(\Lambda^{4}(4\Lambda^{2}+q^{2})^{2}\ q\sqrt{4\Lambda^{2}+q^{2}});$
@@
$\phi_{CN}=-\ln{b_{sl}q^{2}/2.}+\gamma+\ln(1+8/(b_{sl}\Lambda^{2}))-\phi_{1-3})$
@@ V.2 The hadron scattering amplitude
At small momentum transfer in the CNI region, there are two contributions coming from the electromagnetic and the strong interactions.
On the one hand, the interference of such contribution gives the possibility to determine the size of the real part of the scattering amplitude. On the other hand, to determine the form of the imaginary part of the hadronic amplitude, it is necessary to extract the Coulomb contribution and the interference term. As the electromagnetic amplitude and its main contribution as $t\rightarrow 0$ are well known, the measure of experimental data at very small $t$ gives the possibility to improve the normalization of the differential cross sections.
The existence of the different sets of experimental data also allows one to improve separate normalization of experimental data. To compare different sets, some independent basis is needed.
Let us take the calculations of the differential cross sections carried out in the framework of the new high energy generalized structure (HEGS) model [33, 34] as such a basis. The model has only a few free parameters and it quantitatively describes experimental data in a wide domain of the momentum transfer, including the data in the CNI region, in a very wide energy region (from $\sqrt{(}s)=6$ GeV up to LHC energies) simultaneously with the same numbering of the free parameters. The HEGS model assumes Born terms for the scattering amplitude which get unitarization via the standard eikonal representation to obtain the full elastic scattering amplitude.
The scattering amplitude has exact $s\leftrightarrow u$ crossing symmetry as it is written in terms of the complexified Mandelstam variable $\hat{s}=se^{-i\pi/2}$ that determines its real part. The scattering amplitude also satisfies the integral dispersion relation at large $s$. It can be thought of as the simplest unified analytic function of its kinematic variables connecting different reaction channels without additional terms for separate regions of momentum transfer or energy. Note that the model reproduces the diffraction minimum of the differential cross section in a wide energy region [80]. The HEGS model describes the experimental data at low momentum transfer, including the Coulomb-hadron interference region, and hence includes all five electromagnetic spin amplitudes and the Coulomb-hadron interference phase.
@@
$F_{Pom2}^{Born}(\hat{s},t)=$ $h_{Pom2}\ G^{2}_{em}(t)F_{a}(\hat{s},t);$ (33)
$F_{Pom3}^{Born}(\hat{s},t)=$ $h_{Pom3}\ A^{2}_{gr}(t)F_{b}(\hat{s},t);$
$F_{Odd3}^{Born}(\hat{s},t)=$ $h_{Odd3}\ A^{2}_{gr}(t)F_{b}(\hat{s},t);$
where $F_{a}(\hat{s},t)$ and $F_{b}(\hat{s},t)$ have the standard Regge form:
$F_{a}(s,t)\ =\hat{s}^{\alpha_{0}+\alpha_{1}t};F_{b,c}(s,t)\ =\hat{s}^{\alpha_{0}+\alpha_{2}t},$ (34)
with $\hat{s}=s\ e^{-i\pi/2}/s_{0}$ ; $s_{0}=4m^{2}_{p}\ {\rm GeV^{2}}$. The intercept of all main parts of the scattering amplitudes $1.111+\alpha_{0}=$ 1.11 was chosen as arithmetic means from different works including those on inelastic scattering. Hence, at the asymptotic energy we have the universality of the energy behavior of the elastic hadron scattering amplitudes. The main part of the slope of the scattering amplitude has the standard logarithmic dependence on the energy $B(s)=\alpha_{1,2}\ ln(\hat{s})$ with $\alpha_{1}=0.24$ GeV-2 and $\alpha_{2}=\alpha_{1}/4$ GeV-2. It is taken with some correction (eq. 15).
@@
where both (electromagnetic and gravitomagnetic) form factors are used. The corresponding parameters are determined by the fitting procedure.
Figure 5: The HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=13$ TeV (points - experimental data of the TOTEM [27] Collaborations).
Figure 6: The comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=13$ TeV (experimental data [60], $p\bar{p}$ scattering at $\sqrt{s}=53$ GeV (experimental data [61]) and $pn$ scattering at $\sqrt{s}=23$ GeV (experimental data [61]).
The final elastic hadron scattering amplitude is obtained after unitarization of the Born term [83]. There are some different approaches for summing the ladder diagrams - K-marix, U-matrix, standard eikonal, eikonal ( with some inelastic corrections [85]). We analysed these forms and compared the results given by these approaches in some our works (for example, [82, 83, 84] In most part, at high energy ( especially at LHC energy) other approaches give worse results. Moreover, we introduced some parameters reflecting the Kaidalov corrections and obtained that their sizes are near zero or unity. So, this is conserves the standard eikonal form. In a recent work [86] the authors compare the U-matrix with the eikonal and write in the conclusion ”As it is seen in Fig.3 the eikonal approach better agrees with the ATLAS 13 TeV data.”
So, we use the eikonal representation and at first, we have to calculate the eikonal phase
@@
Figure 7: The comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering in the dip-bump regions at $\sqrt{s}=7$ TeV (dashed line and open triangles down), $\sqrt{s}=53.8$ GeV (solid line and circles) and at $\sqrt{s}=19.4$ GeV ( tiny dashed line and boxes).
Figure 8: The comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=27.4$ GeV [left];
and $p\bar{p}$ scattering at $\sqrt{s}=30$ GeV [right].
So to extend the model to low energies, it is necessary to take into account the contributions of the second Reggions. To avoid a substantially increasing number of fitting parameters, we introduce the effective terms which represent the contributions of the sums of different Reggions. For the terms with the energy dependence order $1/\sqrt{\hat{s}}$ we take for the $pp$ scattering the cross even part in the form
$F_{R1}=ih_{R1}/\sqrt{\hat{s}}e^{b_{R1}t\ln{\hat{s}}};$ (36)
and for the $p\bar{p}$ the cross odd term in the form
$F_{R1}=h_{R1}/\sqrt{\hat{s}}e^{b_{R1}t\ln{\hat{s}}}$ (37)
where the value $b_{R1}$ is fixed by unity. For the cross even part fast decreasing with growing energy, we take a term in the form
$F_{R2}=h_{R2}/\hat{s}e^{b_{R2}t\ln{\hat{s}}};$ (38)
The model is very simple from the viewpoint of the number of terms of the scattering amplitude and fitting parameters. There are no any artificial functions or any cuts which bound the separate parts of the amplitude by some region of momentum transfer. In the framework of the model, the description of experimental data was obtained simultaneously at the large momentum transfer and in the Coulomb-hadron region in the energy region from $\sqrt{s}=6$ GeV up to LHC energies. The model gives a very good quantitative description of the recent experimental data at $\sqrt{s}=13$ TeV in the region of the diffraction minimum [31] (see Fig. 5).
As a result, good descriptions of the experimental data of $pp$ and $p\bar{p}$ elastic scattering (90 experimental sets including 4326 points and 256 polarization data) were obtained. Moreover a good description of $pn$ data was also obtained. Some examples of such descriptions are presented in Fig. 5-7. Of particular interest is Fig. 6 where the comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=13$ TeV, $p\bar{p}$ scattering at $\sqrt{s}=53$ GeV, and $pn$ scattering at $\sqrt{s}=23$ GeV at small momentum transfer is shown. Practically for the first time, a simultaneous research of proton-proton, proton-antiproton and proton-neutron elastic scattering has been carried out in a wide energy (from $3.6$ GeV up to $13$ TeV) and momentum transfer region (from $|t|=2.10^{-4}$ GeV2 up to $|t|=14$ GeV2).
In the fitting procedure based on the modern version of FUMILY [87, 88], we used only statistical errors. Systematic errors, which are mostly determined by indefiniteness of luminosity, were taken into account as an additional normalization coefficient, which is the same for all the data of a given set. The different normalization coefficients have practically random distributions. As a result, a wide range of possible forms of the scattering amplitudes prettily decreases.
Figure 9: The comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=19.45$ GeV [left];
and at $\sqrt{s}=12.3$ GeV [right].
Figure 10: The comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=9.9$ GeV [left];
and at $\sqrt{s}=6.1$ GeV [right].
A simultaneous description of the cross sections and spin correlation parameter of different nucleon-nucleon reactions, including 90 sets of experimental data, with the total number of data $N=4326$ gives a very reasonable $\sum_{i,j}\chi^{2}_{i,j}=4826$. The $pn$ case with $526$ experimental data, where the basic parameters were fixed from $pp$ and $p\bar{p}$ scattering, $\sum_{i,j}\chi^{2}_{i,j}=585$.
Figure 12: $\sigma_{tot}(s)$: dashed line - $p\bar{p}$ scattering, solid line -$pp$ scattering; the same but without the second Rejions contributions (tiny dashed - $p\bar{p}$ and long dashed $pp$ scattering).
@@
The data of the TOTEM at $\sqrt{s}=7$ TeV are consistent and their mean value is equal to $98.5$ mb. The ATLAS Collaboration, using their differential cross section data in a region of $t$ where the Coulomb-hadron interference is negligible, obtained the value $\sigma_{tot}=95.35\pm 2.0$ mb. The difference between the two results, $\sigma_{tot}(s)$(T.) - $\sigma_{tot}(s)$(A.) = $3.15$ mb, is about 1 $\sigma$. At $\sqrt{s}=8$ TeV, the measured value of $\sigma_{tot}$ grows, especially in the case of the TOTEM Collaboration, and the difference between the results of the two collaborations grows to $\Delta(\sigma_{tot}(s)$(T.) - $\sigma_{tot}(s)$(A.) $=5.6$ mb, i.e. 1.9 $\sigma$. This is reminiscent of the old situation with the measurement of the total cross sections at the Tevatron at $\sqrt{s}=1.8$ TeV via the luminosity-independent method by different collaborations.
Of course, we cannot say that the normalization of the ATLAS data is better than that of the TOTEM data simply because it coincides with the HEGS predictions. But this exercise may point to the main reason for the different values of the total cross sections obtained by the two collaborations. This does not exclude some further problems with the analysis of experimental data, e.g. those related to the analysis of the TOTEM data at $\sqrt{s}=7$ TeV [89].
One of the most important properties of the diffraction scattering is the form and energy dependence of the dip-bump structure. In Fig. 7, our model calculations are compared with the experimental data of $pp$ scattering at ISR energies ( $\sqrt{s}=19.4$ GeV and $\sqrt{s}=52.8$ GeV) and at the LHC energy $\sqrt{s}=7$ TeV and $\sqrt{s}=13$ TeV. At $\sqrt{s}=19.4$ GeV the real part of the $pp$ scattering changes its sign at $t=0$ and we can see that it gives a small contribution to the diffraction minimum too. It is interesting that at $\sqrt{s}=13$ TeV the diffraction minimum has a sharp form either. The ratio of the maximum to minimum of the differential cross section at $\sqrt{s}=19.4$ GeV is $2.4$ and equals $1.8$ at the LHC energy. The position of the diffraction minimum $t_{min}(s,t)$ moves to low momentum transfer continuously [80] and is described by the model very well (see, for example, Fig. 7 and Figs. 8,9,10 at low energies). It is interesting that the velocity of changing the position of the diffraction minimum changes very slowly. For example, from ISR energy $\sqrt{s}=53$ GeV up to SPS energy $\sqrt{s}=540$ GeV this position changes with a speed of $0.11$ GeV2 per $100$ GeV. Between $\sqrt{s}=540$ GeV and $\sqrt{s}=7$ TeV such speed is two times less and equals $0.006$, at last the position of minimum between $7$ and $13$ TeV changes with a speed of $0.002$ per $100$ GeV. The scaling of this process can approximately be represented as $t_{min}\ln{s/s_{0}}=const$. After the second bump the slope of the differential cross sections increases with energy. It corresponds to the grows of the slope of the diffraction peak. In Figs. 7-9 the model calculations are compared with experimental data at low energies at small momentum transfer for $pp$ and $p\bar{p}$ scattering. It can be seen that the model reproduce well the existing experimental data of both reactions in that region.
@@
The point of $t$, where the imaginary part changes its sign, determines the position of the diffraction minimum. But it moves slightly at some large $t$ by the contribution of the real part of the elastic hadron scattering amplitude. The behavior of the imaginary part of the scattering amplitude over momentum transfer calculated in the framework of the HEGS model is presented in Fig. 11 for the energies $\sqrt{s}=0.51,7,13$ TeV. Again, we can see the point of crossover in the region of $|t|=0.2$ GeV2. Despite the essential growth of the size of the imaginary part of the scattering amplitude at a very small momentum transfer, its slope slightly changes with $t$ in the region of the Coulomb nuclear interference. The size of the slope is practically proportional to the size of the total cross section in that region. However at larger $t$, for example at $|t|=0.1$ GeV2, it grows essentially faster.
@@
It should be noted that the size of the slope of the differential cross sections is determined in that region ofmomentum transfer by the CNI interference term which is proportional to $\alpha_{em}/t$ ( where $\alpha_{em}$ is a fine electromagnetic structure constant). It allows us to analyze [90] the first points of the unique experiment carried out by the ATLAS Collaboration [91].
The $s$ dependence of the total cross sections of $pp$ (solid line) and $p\bar{p}$ (dashed line) scattering is shown in Fig. 12. The difference of the $\sigma^{p\bar{p}}_{tot}$ and $\sigma^{pp}_{tot}$ is less than $1$ mb at $\sqrt{s}=20$ TeV.